Contents
- 1 What is the probability that at least two of them celebrate their birthday on the same day of the year?
- 2 What is the probability that at least one of them celebrates their birthday on the same day as you?
- 3 Was there a day when no one was born?
- 4 How is the birthday paradox used in cryptography?
- 5 How to calculate the probability of a birthday?
What is the probability that at least two of them celebrate their birthday on the same day of the year?
Therefore, the probability that at least 2 will have the SAME birthday is 1- 0.9917958341152186 = 0.008204 .
What is the probability of having the same birthday as someone else?
You can test it and see mathematical probability in action! The birthday paradox, also known as the birthday problem, states that in a random group of 23 people, there is about a 50 percent chance that two people have the same birthday.
What is the probability that at least one of them celebrates their birthday on the same day as you?
The probability of being born any day of the year is 1 or more specifically: 365/365. Since Person B must be born on the same day as Person A their probability is 1/365.
What is the formula for the birthday paradox?
So the chance that two people don’t share a birthday is (365×364)/365². Subtract that from 1 and you get what you expect: that there’s a 1 in 365 chance that two people share a birthday. I could have used factorials to express the numerator, but this way works better.
Was there a day when no one was born?
December 6th is a special day at Who2: it’s the only day of the year on which nobody in our database was born. That’s 2843 famous people (and counting) and none of them born on December 6th.
How is the birthday paradox related to match?
The same principle applies for birthdays. Instead of finding all the ways we match, find the chance that everyone is different, the “problem scenario”. We then take the opposite probability and get the chance of a match. It may be 1 match, or 2, or 20, but somebody matched, which is what we need to find.
How is the birthday paradox used in cryptography?
Here are a few lessons from the birthday paradox: n is roughly the number you need to have a 50% chance of a match with n items. 365 is about 20. This comes into play in cryptography for the birthday attack. Even though there are 2 128 (1e38) GUID s, we only have 2 64 (1e19) to use up before a 50% chance of collision.
How to calculate the reverse formula for the birthday paradox?
To apply this approximation to the first expression derived for p (different), set x = -a / 365. Thus, By putting the value of 1 – a/365 as e -a/365, we get following. By taking Log on both sides, we get the reverse formula.
How to calculate the probability of a birthday?
The latter probability is computed in the following example. 1 365. . (This does not match the real world, as certain birthdays are more common than others, but it is a useful simplification.)